OpenStudy (anonymous):

what is the surface area of a sphere that has a diameter of 20 cm?

Directrix (directrix):

Surface Area of a Sphere = 4 pi r^2 SA = 4 * pi* 10^2 SA = ? @spaghettigoo123

OpenStudy (anonymous):

1256?

jimthompson5910 (jim_thompson5910):

if you are using 3.14 for pi, you are correct

OpenStudy (anonymous):

that's what I did

jimthompson5910 (jim_thompson5910):

ok thought so, you got it correct

OpenStudy (anonymous):

is that all I do?

jimthompson5910 (jim_thompson5910):

yep, plug in r = 10 (diameter = 20 ---> radius = 10) and evaluate

OpenStudy (anonymous):

oh I see

OpenStudy (anonymous):

thanks guys!

jimthompson5910 (jim_thompson5910):

yw